Monthly Archives: August 2014

The week of August 25

The software group continues to test different aspects of the TCS. They found many of the bugs that stopped us from generating trajectories last week. We should now be able to generate trajectories at any Az and for normal sidereal objects and for geostationary satellites (useful for diagnostics). We do
continue to run into some small teathing problems with the new tracker as we start running it through its paces, e.g. loose connectors, cables that are not routed the correct way and blown fuses.

The electrical and mechanical teams worked in the Krabbenhut this week and was able to remove the old Champion compressor, install new vertical air tank and move the Power-X compressor which supplies the instrument air. This was all done to make room for the Nitrogen Vaporizor which will generate very pure N2 gas for instrument purging from our large LN tank.

Our electronics tech continues to work with the Ultra-pure water system trying to add reliability and consistency to the system. He can now produce 17.1 Mohms water (any reading above 10 Mohms is considered is considered Ultra High Purity for water, with 18.2 Mohms being the ultimate maximum).

The mirror team finished the piston work this week and we hope to have the next round of piston be done with the DMI rather than with the sphereometer.

In the Facility Thermal Management Project Veliz Construction completed the welding of pylon frames (to support ducting) together in place over side of the mountain and ARC is working on the ducting next to the Krabbenhut and down the side of the mountain. ARC mechanical has completed the WFU glycol manifold at the outside of the dome. This is the point where HET personnel will continue the piping layout into the dome.

The week of August 18

This week the mirror team began making piston moves of the individual segments that make up the primary mirror. We began with the central mirror, of which we base most mirrors moves, and then moved to those mirrors which had small enough piston moves to be controlled by SAMS. Next week we will move on to those mirrors whose piston will require us to make moves outside of the capture range of SAMS including those mirrors where the piston moves are out of range of the actuators and thus requires moves of the M30 frame screws. These were not completed this week because the mirror team is also interviewing two candidates for the Opto-mechanical position.

Veliz Construction continues placing the pylons to support the K-Hut exhaust duct and fan. They are working on the pylons over the side of the mountain. ARC Mechanical is busy building the HDPE piping run around the control/service building slab for the WFU glycol system. They also began welding the copper pipe in place in the pipe vault for the Mitsubishi compressors.

We have also begun integrating the structure sub-system into requested trajectory generation and execution. Having some problems with satellite and ephemeris trajectories.

Much of the rest of the team was working on preparations for a move of the instrument air compressor inside the Krabbenhut to make room for the Nitrogen Vaporizor. Other work includes installing Anemometers, purge lines for the new tracker, transient suppression for the tracker electronics and other various activities to keep the place running.

The week of August 11

The blogger is back from a summer vacation. Plenty of activity at the telescope…

The 97 point test was completed a few weeks back and that data in under analysis. A mount model based on the previous 21 point test has been installed and tested. In the last 2 weeks a great deal of reliability has been added to the tracker software. One of the mechanical engineers from Austin was out this week reinstalling the tip/tilt camera, DMI and a new sky camera. On this last wednesday they use that camera to acquire a few fields and make sure the basic pointing of the telescope has not been lost. In addition to a few bright stars the group observed an open cluster just to see what the image quality might be for future plate scale analysis. Today the tip-tilt camera and DMI were used during several trajectories. This is a small milestone. We can use this data to refine the mount model with respect to the primary mirror (best done after we piston the array).

The mirror team has installed the last 4 mirrors into the array and the array is fully populated. We will start mirror swaps when the Remote Thermal Area project is complete and we have a chance to commission the strip and wash room for acid vapor removal. In the mean time the mirror team have measured the array with a sphereometer so that we can implement a piston move in the coming week.

The contractors on site have finished the silver tape application to the control and services building and next week they will start on the Krabbenhut. The mechanical contractor continue to work on the insulation that goes around the lines/pipes out to the remote thermal area. Last week the dirt works contractors finished the piers down the side of the mountain which will carry the acid vapor from the strip and wash room as well as the heat from the control and services building. The contractors are now working on the acid ducting inside the building and the conduit piers on the side of the Krabbenhut. The electrical contractors completed the move of one of the transformers from outside the control and services building to inside the Krabbenhut. This required the HET to operate from a generator for nearly a week. Big thanks to the McDonald Physical Plant team for allowing us to use the generator and keeping it filled over the weekend so that our sewer system would not be in jeopardy.